# Do Students Write Better Post-AI Support?

### v1

@@ 5.5 Limitations and Future Directions
Several limitations warrant attention when interpreting the present findings. First, the current study assessed students’ writing performance solely at two temporally adjacent points (with and immediately after the removal of AI scaffolding) within the same experimental session, limiting our ability to make broader claims about delayed or long-term improvements. Future longitudinal studies involving follow-up measures conducted at various time intervals can more definitively establish delayed effects or longer-term skill consolidation. Second, the task’s particular context, healthcare simulations involving multimodal academic writing and visual data interpretation, limits direct generalisability across disciplines or other academic writing types. Replicating or expanding this approach to diverse fields (such as STEM, social sciences, and the humanities) would clarify the robustness and generalisability of current findings. Third, although we distinguished passive from proactive chatbot conditions, additional chatbot features (e.g., specific scaffolding strategies, transparency, adaptivity, fairness) likely shape interactions in complex ways, warranting follow-up studies that evaluate nuanced chatbot design variables. Fourth, qualitative evidence clarifying students’ cognitive processes, metacognitive reflections, and explicit interactions with GenAI tools was beyond our current analysis scope. Future research integrating qualitative approaches (think-aloud protocols, interviews, reflection statements) or detailed log-file analyses of user-AI interactions could richly explain underlying adjustments or strategic variations driven by differing GenAI literacy levels.
@@ 6 Conclusion
As GenAI technologies like advanced chatbots become prevalent in education, questions arise regarding their impact on independent thought and student performance. Our study provides empirical evidence that GenAI literacy positively predicts students’ multimodal academic writing performance even after the removal of AI support. Significantly, these GenAI literacy skills influenced outcomes beyond domain-specific knowledge and prior academic performance, highlighting their distinct importance. Additionally, interaction style mattered; passive chatbot support effectively promoted autonomous skill application among AI-literate learners. These results underscore the importance of explicitly nurturing GenAI literacy within curricula and thoughtfully balancing AI assistance with learner autonomy. Future longitudinal studies should explore lasting impacts, informing the responsible integration of GenAI into educational settings.
*Acknowledgments This research was in part supported by the Australian Research Council (DP220101209, DP240100069). L.Y.’s work is fully funded by the Digital Health CRC (Cooperative Research Centre). D.G.’s work was, in part, supported by the DHCRC and Defense Advanced Research Projects Agency (DARPA) through the Knowledge Management at Speed and Scale (KMASS) program (HR0011-22-2-0047). The DHCRC is established and supported under the Australian Government’s Cooperative Research Centres Program. The U.S. Government is authorised to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright notation thereon. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of DARPA or the U.S. Government.
*Conflict of interest
The authors declare no potential conflict of interest.
*Ethics Statement Ethics approval was obtained from Monash University (Project ID: 37307). Informed written consent was also obtained from students before the data collection.
*Data Availability
For any queries regarding the data and analysis codes, please contact the corresponding author.

### v2

@@ 5.6 Limitations and Future Directions
Several limitations warrant attention when interpreting the present findings. First, the current study assessed students’ writing performance solely at two temporally adjacent points (with and immediately after the removal of AI scaffolding) within the same experimental session, limiting our ability to make broader claims about delayed or long-term improvements. Future longitudinal studies involving follow-up measures conducted at various time intervals can more definitively establish delayed effects or longer-term skill consolidation. Second, the task’s particular context, healthcare simulations involving multimodal academic writing and visual data interpretation, limits direct generalisability across disciplines or other academic writing types. Replicating or expanding this approach to diverse fields (such as STEM, social sciences, and the humanities) would clarify the robustness and generalisability of current findings. Third, although we distinguished reactive from proactive chatbot conditions, additional chatbot features (e.g., specific scaffolding strategies, transparency, adaptivity, fairness) likely shape interactions in complex ways, warranting follow-up studies that evaluate nuanced chatbot design variables. Fourth, qualitative evidence clarifying students’ cognitive processes, metacognitive reflections, and explicit interactions with GenAI tools was beyond our current analysis scope. Future research integrating qualitative approaches (think-aloud protocols, interviews, reflection statements) or detailed log-file analyses of user-AI interactions could richly explain underlying adjustments or strategic variations driven by differing GenAI literacy levels. Finally, we did not directly measure several related literacies (e.g., information/digital literacy, baseline academic writing proficiency). Although our controls (VLAT, AI-support scores) and robustness checks reduce the risk that GLAT effects reflect a proxy for other literacies, residual confounding cannot be fully ruled out. Future work should adopt multitrait–multimethod designs (e.g., combining GLAT, academic writing assessments, and digital/information literacy scales) and incorporate language proficiency screens to strengthen discriminant validity.
@@ 6 Conclusion
This paper examined how GenAI literacy shapes students’ multimodal academic writing with and without AI support. We conducted an experimental study with 79 higher education students who wrote with either reactive or proactive chatbot assistance followed by an independent task without AI support. Results showed that GenAI literacy predicted independent writing only in the reactive condition, where students had to self-initiate and regulate their engagement, but not in the proactive condition, where scaffolding equalised outcomes. Regression confirmed that this effect was robust beyond domain knowledge and prior AI-assisted performance, and mediation analysis revealed no indirect pathway, highlighting a direct influence of GenAI literacy. We introduce the concept of the agency gap to capture this pattern and argue that GenAI literacy is a distinct, task-general competence whose relevance depends on interaction design. The study contributes theoretically by clarifying when and how GenAI literacy matters, and practically by identifying two levers for equitable design: cultivating literacy to foster independence and deploying scaffolds to mitigate disparities.
Acknowledgments
Hidden for review purposes.
Data availability
Data will be made available on request.
Appendix A Generative AI Literacy Assessment Test (GLAT)
| Item | Dimension | Question | Options (Answer Bolded) |
| 1 | Know & Understand | Which of the following best describes "Generative AI"? |
A. AI that creates new content like text, images, or music by learning from existing data.
B. An AI system designed to enhance the speed and accuracy of data retrieval in search engines. C. A form of artificial intelligence that focuses on translating languages in real-time. D. AI technology used primarily for managing and organizing large databases. |
| 2 | Know & Understand | Which of the following statements best describes an LLM (Large Language Model)? |
A. It generates text by analyzing and summarizing large volumes of web content.
B. It generates text by predicting the next word based on the context of previous words. C. It generates text by translating input text into multiple languages simultaneously. D. It generates text by using pre-defined templates and filling in the blanks. |
| 3 | Know & Understand | Which of the following tasks can Generative AI perform with a high degree of accuracy? |
A. Predicting stock market trends
B. Making ethical decisions in complex scenarios C. Diagnosing rare diseases D. Generating human-like text based on prompts |
| 4 | Know & Understand | In the context of Generative AI, what is "zero-shot learning"? |
A. Training a model without any data.
B. The ability of a model to perform a task without any task-specific training. C. A method of reducing the model’s training time to zero. D. A technique for generating synthetic training data. |
| 5 | Know & Understand | Which of the following is a potential challenge when using prompt-based development for text generation? |
A. The language model can only generate binary outputs.
B. The need for extensive labeled data to train the model. C. Crafting a prompt that accurately captures the desired context and nuances. D. The requirement for complex feature engineering. |
| 6 | Know & Understand | What does the term "token" refer to in the context of a large language model (LLM)? |
A. A token is a unit of text, such as a word or a subword, that the model processes individually.
B. A token is a unique identifier assigned to each user interacting with the language model. C. A token is a security measure used to authenticate API requests to the language model. D. A token is a reward given to users for contributing valuable data to train the language model. |
| 7 | Know & Understand | Which of the following is NOT a requirement for an AI to be considered artificial general intelligence (AGI)? |
A. The ability to learn and adapt to new tasks without human intervention.
B. The capability to perform tasks across various domains with human-like proficiency. C. The ability to predict future events with perfect accuracy. D. The capacity to understand and generate natural language. |
| 8 | Know & Understand | How does RAG (Retrieval-Augmented Generation) enhance the capabilities of an LLM? |
A. By improving its grammar and syntax.
B. By providing it with real-time and relevant data. C. By increasing its computational speed. D. By enabling it to understand multiple languages. |
| 9 | Use & Apply | When using generative AI to create a marketing pitch, which of the following strategies is least likely to be effective? |
A. Supplying the AI with information about the target audience
B. Asking the AI to include unique selling points and benefits C. Requesting the AI to use persuasive language techniques D. Providing the AI with a list of competitors’ products |
| 10 | Use & Apply | After deploying a customer service chatbot, you notice that it frequently provides outdated information about company policies. What is the best course of action to address this issue? |
A. Implement a feedback loop where users can flag outdated information for review.
B. Schedule regular updates to the chatbot’s training data to include the latest company policies. C. Set up a system where complex or policy-related queries are escalated to human agents for accurate responses. D. Conduct a comprehensive audit of the chatbot’s performance metrics to identify areas for improvement. |
| 11 | Use & Apply | Suppose you have a large dataset of emails and you want to build an application to answer questions based on this dataset. Which of the following scenarios best illustrates the advantage of using RAG over prompting (i.e., without RAG)? |
A. You need to generate creative writing pieces based on the email content.
B. You want to ensure the model can answer questions even if it has never seen similar questions before. C. You need to answer questions that require specific information from different parts of the email dataset. D. You want to reduce the size of the language model to save computational resources. |
| 12 | Evaluate & Create | As a student using a Large Language Model (LLM) to gather information for an assignment, how should you approach the information it provides? |
A. The LLM’s answers are always more trustworthy than any information you will find on the internet, so you can use them without further verification.
B. The LLM’s answers are generally more trustworthy than internet sources, but you should still verify the information with other reliable sources. C. The LLM’s answers are not necessarily more trustworthy than internet sources, and you should cross-check the information with other credible references. D. The LLM’s answers are less trustworthy than internet sources because it relies on outdated information. |
| 13 | Evaluate & Create | It is unlikely for an LLM to provide an accurate summary of the latest financial market trends in real-time. Is this statement true or false? |
A. True, because the LLM’s data may be outdated due to its knowledge cutoff.
B. True, because the LLM is not good at handling numbers and structured data. C. False, because the LLM frequently updates its knowledge base. D. False, because the LLM is capable of synthesizing the latest market data automatically. |
| 14 | Evaluate & Create | A generative AI tool has provided a summary of a research paper. The summary states, "The study found that increased screen time is directly correlated with decreased attention spans in children aged 8-12." What is your next step? |
A. Accept the summary as accurate because AI tools are generally reliable.
B. Ask the AI to provide more details about the study’s methodology and results. C. Cross-check the summary with the original research paper. D. Use another AI tool to generate a summary for comparison and evaluate the consistency between both summaries |
| 15 | Evaluate & Create | While reviewing a video of a well-known public figure making controversial statements, which characteristic confirms the video was NOT generated by AI? |
A. The public figure’s voice sounds like themselves.
B. The video has a professional and polished appearance. C. The video is high-quality with smooth transitions. D. None of the above. |
| 16 | Ethics | When a generative AI system is used for screening job applications, what issue might arise concerning the quality and fairness of hiring decisions? |
A. The AI system might overlook applicants’ unique achievements and extracurricular activities.
B. The AI system could misinterpret minor formatting differences in resumes. C. The AI system might not effectively handle applications submitted in various languages. D. The AI system could reinforce existing biases found in historical hiring data. |
| 17 | Ethics | In a healthcare startup, an accurate AI model recommends treatments, but doctors don’t trust it because they can’t understand how the model arrived at its conclusions. What core issue does this scenario illustrate? |
A. The AI model uses obsolete training data.
B. The training dataset lacks sufficient diversity. C. The treatment guidelines input are incorrect. D. The AI model behaves as a black box. |
| 18 | Ethics | What are the potential copyright implications for a journalist using an AI-generated image in a commercial article? |
A. The journalist needs to check the licensing policy of the AI tool they used.
B. The AI-generated image is automatically free to use without any restrictions. C. The journalist must pay a standard licensing fee to use the AI-generated image. D. The image cannot be used in any commercial context because it is AI-generated. |
| 19 | Ethics | Should we impose restrictions on the outputs of generative AI technologies? |
A. Yes, to reduce the computational resources required for operating these technologies.
B. Yes, to prevent the dissemination of harmful or misleading content. C. No, as it would hinder technological innovation and creativity. D. No, because users should have the freedom to access all generated content. |
| 20 | Ethics | Sending personal information to cloud-based generative AI tools has little privacy concerns. |
A. True, as this information is encrypted using sophisticated algorithms during the transmission process.
B. True, as generative AI tools are black-box systems and cannot output personal information even if it is used for model training. C. False, as generative AI tools train on unencrypted data and can output private information based on their probabilistic nature. D. False, as advancements in quantum computing can easily decipher the encrypted data. |
Appendix B Mini-Visualization Literacy Assessment Test (VLAT)
| Item ID | Visualization Type | Question |
| 1 | Line Chart | What was the price of a barrel of oil in February 2020? |
| 2 | Bar Chart | What is the average internet speed in Japan? |
| 3 | Stacked Bar Chart | What is the cost of peanuts in Seoul? |
| 4 | 100% Stacked Bar Chart | Which country has the lowest proportion of Gold medals? |
| 5 | Pie Chart | What is the approximate global smartphone market share of Samsung? |
| 6 | Histogram | What distance have customers traveled in the taxi the most? |
| 7 | Scatterplot | There is a negative linear relationship between the height and the weight of the 85 males. |
| 8 | Area Chart | What was the average price of pound of coffee beans in October 2019? |
| 9 | Stacked Area Chart | What was the ratio of girls named “Isla” to girls named “Amelia” in 2012 in the UK? |
| 10 | Bubble Chart | Which city’s metro system has the largest number of stations? |
| 11 | Choropleth Map | In 2020, the unemployment rate for Washington (WA) was higher than that of Wisconsin (WI). |
| 12 | Treemap | eBay is nested in the Software category. |
Appendix C Rubric for Evaluating Academic Writing with Visual Data
| Metric | Score | Description and Example |
| Insightfulness | Level 5 (Excellent) |
- Provides highly original and profound insights.
- Provides a holistic response that addresses every key element of the analysis. - Draws thought-provoking, actionable conclusions that extend beyond obvious interpretations. Example: "Nurse 1’s disproportionate focus on Amy (50%) coupled with their central communication role and elevated stress levels (111 BPM near Bed 4) indicates an unsustainable workload distribution. Implementing a more balanced task allocation could enhance team efficiency and reduce individual stress." |
| Level 4 (Very Good) |
- Offers detailed and valid analysis with some synthesis of data.
- Provides a well-rounded response but may lack depth in one or two areas or overlook minor elements. - Conclusions are solid but may lack the depth or originality of Level 5. Example: "While focusing on Amy aligns with priorities, Nurse 1’s increased stress due to communication demands suggests that sharing responsibilities might alleviate stress and improve outcomes." |
| Level 3 (Good) |
- Provides accurate analysis based on direct observations.
- Limited synthesis of data; basic, isolated analysis with little connections. - Insights are correct but straightforward and may be somewhat obvious. Example: "Nurse 1 spent more time attending to Amy, which led to higher stress levels compared to Nurse 2." |
| Level 2 (Fair) |
- Makes surface-level observations with minimal analysis.
- Omits key elements or provides superficial explanations. - Demonstrates a limited understanding of the scenario. Example: "Nurse 1 worked on Amy for half the time and had a higher heart rate." |
| Level 1 (Poor) |
- Provides an incomplete or highly superficial response.
- Lacks any meaningful interpretation or insight. Example: "Nurse 1 was stressed with a higher heart rate while spending time on Amy." |
| Visual Data Integration | Level 5 (Excellent) |
- Fully integrates all relevant visual data types into the analysis.
- Draws meaningful connections between different data sources. - Builds a cohesive narrative that supports complex conclusions. Example: "The solo focus (50%) on Amy, Nurse 1’s central role in communication, and elevated heart rate (111 BPM at Bed 4) collectively highlight an imbalance in task distribution, emphasising the need for better teamwork." |
| Level 4 (Very Good) |
- Effectively incorporates multiple types of visual data.
- Makes connections between data types, though some integrations may be less thorough. - Uses visual data to support analysis, but may not fully exploit all available information. Example: "Nurse 1’s extensive time with Amy and higher stress levels near Bed 4 suggest that task allocation affected their stress, as shown in the charts and heart rate data." |
| Level 3 (Good) |
- Utilises at least one type of visual data effectively.
- References other data types but with limited integration. - Analysis supported by visual data but lacks depth in synthesis. Example: "According to the communication network, Nurse 1 talked more, which might explain their higher stress level." |
| Level 2 (Fair) |
- Mentions visual data superficially without meaningful integration.
- Relies on single data points without connecting them to the broader context. Example: "The chart shows Nurse 1 had a heart rate of 111 BPM near Amy." |
| Level 1 (Poor) |
- Fails to incorporate visual data into the analysis.
- May misinterpret or ignore available data. Example: "Nurse 1 was stressed, as seen in the data." |
| Organisation and Coherence | Level 5 (Excellent) |
- Presents ideas in a clear, logical order with excellent structure.
- Uses seamless transitions between points and paragraphs. - Arguments are well-developed and interconnected. Example: "By allocating her solo efforts (50%) to Amy, Nurse 1 became central in patient care. This, combined with her elevated stress levels (111 BPM near Bed 4) and primary communication role, indicates a need for task redistribution to promote balance and reduce burnout." |
| Level 4 (Very Good) |
- Maintains a clear and logical structure with minor lapses.
- Uses transitions effectively, though some areas may be less smooth. - Ideas are well-organised and mostly easy to follow. Example: "Nurse 1’s higher stress, as indicated by heart rate data, relates to her focus on Amy. Improving teamwork could help mitigate this issue." |
| Level 3 (Good) |
- Displays basic organisation but may have occasional clarity issues.
- Transitions between ideas may be abrupt or inconsistent. - Arguments are understandable but may lack depth. Example: "Nurse 1 worked more on Amy and had higher stress. Teamwork might have helped reduce her stress." |
| Level 2 (Fair) |
- Ideas are poorly organised or disjointed.
- Lacks clear transitions, making it hard to follow the argument. - Writing may be repetitive or off-topic. Example: "They worked alone a lot. Nurse 1’s stress was high. Communication is important." |
| Level 1 (Poor) |
- Lacks coherent structure; ideas appear random or jumbled.
- Writing is confusing and difficult to understand. Example: "Nurse work on Amy. Stress high. Data shows tasks." |
| Linguistic Quality | Level 5 (Excellent) |
- Writing is fluent with varied and precise vocabulary.
- Virtually free of grammatical errors. - Complex ideas are expressed clearly and elegantly. Example: "The disproportionate workload on Nurse 1, evidenced by her solo focus on Amy and central communication role, resulted in elevated stress levels. A more equitable task distribution could mitigate stress and improve patient care." |
| Level 4 (Very Good) |
- Generally clear and correct writing with minor errors.
- Vocabulary is appropriate but may lack variety. - Ideas are communicated effectively. Example: "Nurse 1’s higher stress relates to spending more time on Amy and communicating more. Sharing tasks might reduce her stress." |
| Level 3 (Good) |
- Writing is understandable but contains noticeable errors.
- Vocabulary and sentence structure are simple or repetitive. - Errors do not significantly impede understanding. Example: "Nurse 1 worked more with Amy and was more stressed than Nurse 2." |
| Level 2 (Fair) |
- Frequent grammatical errors that may hinder comprehension.
- Limited vocabulary; word choices may be inappropriate or incorrect. - Ideas are difficult to follow due to language issues. Example: "Stress Nurse 1 shows from Amy is cause by talks more and heart rate." |
| Level 1 (Poor) |
- Writing is riddled with errors; meaning is often obscured.
- Extremely limited vocabulary and flawed sentence structure. - Ideas are not effectively communicated. Example: "Nurse stress high. Work Amy. Data show." |
| Critical Thinking | Level 5 (Excellent) |
- Demonstrates deep analysis and understanding of complexities.
- Addresses possible counterarguments or alternative interpretations. - Provides well-reasoned, innovative solutions or recommendations. Example: "While prioritising Amy is essential, Nurse 1’s elevated stress (111 BPM) and communication burden indicate an imbalance. Introducing a shared workload model could distribute tasks more evenly, potentially improving team performance and reducing individual stress." |
| Level 4 (Very Good) |
- Analyses data carefully with awareness of underlying complexities.
- Offers logical conclusions and considers some implications. - Suggestions are practical and grounded in the data. Example: "Nurse 1’s stress suggests that although focusing on Amy is appropriate, collaborating more with colleagues might alleviate stress." |
| Level 3 (Good) |
- Provides basic analysis but does not delve into complexities.
- Conclusions are valid but may be obvious or lack depth. - Limited consideration of broader implications or solutions. Example: "Because Nurse 1 worked more on Amy, she was more stressed." |
| Level 2 (Fair) |
- Offers superficial analysis with little critical engagement.
- Conclusions may be unsupported or overly simplistic. Example: "Nurse 1 was stressed because she focused on Amy." |
| Level 1 (Poor) |
- Lacks critical analysis; may misinterpret data.
- Conclusions are incorrect or irrelevant. Example: "Nurse 1 stressed due to work on Amy. Nurse 2 worked less." |